OpenStudy (anonymous):

Can you help me? I don't get it. For the following problem, predict the next three numbers. -3,1,0,4,___,___,___.

OpenStudy (saifoo.khan):

is there any typo?

OpenStudy (anonymous):

No, typo.

OpenStudy (anonymous):

-3,1,0,4,3,7,6,10,9,13 -3+4=1(add 4) 1-1=0(subtract 1) 0+4=4(add 4) 4-1=3(subtract 1) not sure :( estudier what do you think ????/

OpenStudy (anonymous):

I think that's probably as good as ur going to get, it's not very obvious.

OpenStudy (anonymous):

Ok, so it's 4,3,7?

OpenStudy (anonymous):

Where u get this question?

OpenStudy (anonymous):

It's an old quiz from a while ago. I have to make corrections before I turn it back in. I got this problem wrong.

OpenStudy (anonymous):

Mahalo nui loa, you guys/people :)

OpenStudy (anonymous):

Mahalo nui loa, you guys/people :)